OpenStudy (anonymous):

Write x+2y=5 as a function. Please tell me how you did it too.

OpenStudy (anonymous):

Firstly, find the value of y in that equation.. Take x on other side and try dividing by 2 both the sides, you will get the value of y..

OpenStudy (anonymous):

Can you show me in characters

OpenStudy (unklerhaukus):

\[x+2y=5\] \[x+2y-x=5-x\] \[2y=5-x\] \[y=\frac{5-x}{2}\]

OpenStudy (anonymous):

It doesn't say that in my book. How would I get this:

OpenStudy (anonymous):